# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ VI Outlook
Clearly, one practical application of this protocol will be to adapt it in a fashion where one is able to protect encoded information up to late times. In the current transmon-based setting, the key impediment towards this is due to the noisy offset charge [73, 74], which has been ignored for simplicity in our analysis. At the freezing point, where the Hamiltonian effectively becomes $\mathcal{H}^{0}_{\widehat{n}}$, the spatio-temporal dependence of the offset charge can play a significant role in the decoherence. Finding alternative routes to mitigate this effect in a realistic superconducting device based architecture remains an exciting open direction.
Acknowledgements.
@@
We thank S.D. Börner, A. Das, V. Fatemi, B. Kobrin, Y. Lensky, R. Moessner, E. Rosenberg, P. Roushan, S. Roy, V. Smelyanskiy, C. Tahan, G. Vidal and B. Ware for a number of useful discussions leading up to this work. RM thanks M. Bukov and S. Bandyopadhyay for many useful suggestions regarding QuSpin PYTHON package [75, 76], and thanks S.D. Börner for providing the Julia code dynamical.jl [72] for solving the semiclassical equations for the transmon arrays. RM is supported by a Fulbright-Nehru Grant No. 2877/FNDR/2023-2024 sponsored by the Bureau of Educational and Cultural Affairs of the United States Department of State. HG is supported by a Wilkins postdoctoral fellowship at Cornell University. DC is supported in part by a New Frontier Grant awarded by the College of Arts and Sciences at Cornell University and by a Sloan research fellowship from the Alfred P. Sloan foundation.
Appendix A Analytical details of the Floquet-Magnus Expansion
@@
In this appendix, we provide additional details regarding the Floquet-Magnus expansion.
A.1 General Expression for the Floquet Hamiltonian
@@
$U_{I}(t)={\mathcal{T}}\exp\left(-i\int_{0}^{t}{\rm d}t^{\prime}\mathcal{H}_{\text{mov}}(t^{\prime})\right)\,,$ (A.1)
@@
$i\partial_{t}U_{I}(t)=\mathcal{H}_{\text{mov}}(t)U_{I}(t)\,.$ (A.2)
@@
$\begin{split}\partial_{t}e^{\Omega(t)}&=\int_{0}^{1}{\rm d}\lambda e^{\lambda\Omega(t)}\partial_{t}\Omega(t)e^{(1-\lambda)\Omega(t)}\\ &=h({\rm ad}_{\Omega(t)})\partial_{t}\Omega e^{\Omega(t)}\,,\end{split}$ (A.3)
where $h(x)=(e^{x}-1)/x$ and ${\rm ad}_{X}Y=[X,Y]$. Combining this with Eq.A.2, we obtain,
$\partial_{t}\Omega(t)=-ih\left({\rm ad}_{\Omega(t)}\right)\mathcal{H}_{\text{mov}}(t)\,.$ (A.4)
@@
$\mathcal{H}_{\text{eff}}^{0}$ $=\frac{1}{T}\int_{0}^{T}{\rm d}t\mathcal{H}_{\text{mov}}(t)\,,$ (A.5)
$\mathcal{H}_{\text{eff}}^{1}$ $=\frac{1}{2iT}\int_{0}^{T}{\rm d}t_{1}\int_{0}^{t_{1}}{\rm d}t_{2}\left[\mathcal{H}_{\text{mov}}(t_{1}),\mathcal{H}_{\text{mov}}(t_{2})\right]\,,$ (A.6)
$\mathcal{H}_{\text{eff}}^{2}$ $=\frac{-1}{6T}\int_{0}^{T}{\rm d}t_{1}\int_{0}^{t_{1}}{\rm d}t_{2}\int_{0}^{t_{2}}{\rm d}t_{3}$ (A.7)
@@
Figure 12: (a) ${\cal{F}}$ (Eq.  16) averaged over 100 Floquet cycles as a function of $\alpha=A/\omega$ for two different system sizes ($L=6,8$) and onsite Hilbert space (${\cal{H}}_{D}=2$), with |$\ket{\psi(0)}$ chosen as ground state of $H(0)$. In addition to the sharp peaks at even integer $\alpha$, peaks at odd integers are also seen. (b) Same as in (a), but with ${\cal{H}}_{D}=3$. With increasing $\alpha$, $\cal{F}$ flattens out since $\mathcal{H}^{0}_{\widehat{\varphi}}$ weakens compared to $\mathcal{H}^{0}_{\widehat{n}}+f(t)\mathcal{H}^{\rm{drive}}_{\widehat{n}}$. (c) For $L=8$ transmons and ${\cal{H}}_{D}=3$, variation of ${\cal{F}}$ for different |$\ket{\psi(0)}$, drawn from the eigenspectrum of $H(0)$ in Fig. 13. All other parameters are as in Fig. 2.
Figure 13: The eigenspectrum of $H(0)$ for the square-wave drive, plotted with the corresponding Fock state indices. The eigenstates marked by arrows are used as |$\ket{\psi(0)}$ to calculate the observables in Eq. 16 and Fig. 12.
A.2 Magnus Expansion of the Transmon Hamiltonian
In this section we explicitly evaluate Eqs.  A.5-A.7 for the transmon Hamiltonian. The co-moving Hamiltonian is given by
$\mathcal{H}_{\text{mov}}(t)=\mathcal{H}^{0}_{\widehat{n}}-E_{J}\sum_{k}\cos({\widehat{\varphi}}_{k}-\theta_{k}(t))\,,$ (A.8)

### v2

@@ Discussion
Clearly, one practical application of this protocol will be to adapt it in a fashion where one is able to protect encoded information up to late times. In the current transmon-based setting, the key impediment towards this is due to the noisy offset charge 70, 71, which has been ignored for simplicity in our analysis. At the freezing point, where the Hamiltonian effectively becomes $\mathcal{H}^{0}_{\widehat{n}}$, the spatio-temporal dependence of the offset charge can play a significant role in the decoherence. Finding alternative routes to mitigate this effect in a realistic superconducting device based architecture remains an exciting open direction.
Methods
Here we introduce the basics of Floquet theory that underpins the theory throughout this work and introduce the correspondence between the quantum and semi-classical limits of the coupled transmon system.
Floquet Theory
Our time-dependent Hamiltonian $H(t)=H_{0}+H_{\rm{D}}(t)$ consists of two parts — the static piece, $H_{0}\equiv{\cal{H}}_{\widehat{n}}^{0}+{\cal{H}}_{\widehat{\varphi}}^{0}$, and the driven piece, $H_{\rm{D}}(t)\equiv f(t)\mathcal{H}^{\rm{drive}}_{\widehat{n}}$. Due to the discrete time translational symmetry associated with the period $T\equiv 2\pi/\omega$, the solutions of the Schrödinger equation can be written as |$\ket{\psi_{\alpha}(t)}=e^{-i\epsilon_{\alpha}t}\ket{u_{\alpha}(t)}$, with time-periodic Floquet wavefunction, |$\ket{u_{\alpha}(t+T)}=\ket{u_{\alpha}(t)}$, and Floquet quasi-energy, $\epsilon_{\alpha}\in(-\omega/2,\omega/2]$, respectively. The Floquet-Schrödinger equation is written as 72, 73,
[…]
In the limit of both large driving amplitude and frequency (relative to the transmon frequency given in Eq. (4)), $H_{\rm eff}$ can be obtained from the modified magnus expansion 33, which we review now. We consider moving into the co-moving frame of the driving term, via the unitary transform
$W(t)=\exp(-i\theta(t)\mathcal{H}_{\widehat{n}}^{\text{drive}})\,,$ (34)
[…]
Here we use the notation $\mathcal{H}^{0}_{\widehat{n}}[\widehat{n}_{i}]$ to indicate that $\mathcal{H}^{0}_{\widehat{n}}$ is being written as a function the operator $\widehat{n}_{i}$ and similarly for other Hamiltonians throughout the paper. The Magnus expansion 33 states that the leading order Floquet effective Hamiltonian is given by the one-period average of $\mathcal{H}_{\text{mov}}(t)$:
For the model under consideration, we find
Here $\dots$ denote higher order corrections in $1/\omega$ which will be studied in further detail in Sec. LABEL:results:FM; see also Supplemental Methods B. Therefore, we find at a set of isolated points, the non-commuting $\varphi$-dependent terms vanish, and the effective Hamiltonian becomes seemingly integrable (at leading order):
${\rm{Case~{}I:}}$ $~{}~{}\frac{A}{\omega}\approx 2n,\ {n\in\mathbb{Z}},$ (40a)
Semi-Classical Limit
The dynamics of the coupled transmons is amenable to a semi-classical analysis, where we replace the non-commuting operators $\widehat{\varphi}$ and $\widehat{n}$ by classical real variables, and treating the commutator bracket as a Poisson bracket, $\{\varphi_{i},n_{j}\}=\delta_{ij}$, respectively. Recall that in the absence of a Floquet drive, the semi-classical dynamics of transmon arrays has been analyzed, and the behavior is generically expected to be chaotic 74, 45 without substantial frequency detuning or tunable coupling 46. Our goal will be to study how the Floquet drive in the absence of any other strategy helps arrest the growth of chaos, and protect the memory associated with generic initial states.
The Hamilton equations of motion for the coupled set of phase and number variables in Eq. (1a) are given by,
where $\sum_{j\in\text{adj}_{i}}$ extends over the nearest neighbors of $i$. In the language of pendula the interaction between the transmons is a momentum-momentum interaction that is due to the capacitive coupling. In addition to obtaining the full dynamics associated with $\varphi_{i}(t),~{}n_{i}(t)$, we will also diagnose chaos by computing the Lyapunov exponent 75. Specifically, we focus on the maximal Lyapunov exponent (MLE), $\lambda_{L}$, which determines the rate at which trajectories that are initially “close” diverge with time, $\delta\pi(t)\approx\delta\pi\exp{(\lambda_{L}t)}$. We define $\lambda_{L}$ as follows 76. Consider two trajectories in phase space with an infinitesimal initial separation $\delta\pi_{0}$ at time $t=0$, and let $\delta\pi(t)$ be the resulting separation at a later time $t$. Then
By contrasting the Lyapunov exponent for the static vs. driven transmon arrays, we highlight in Results LABEL:results:SC how dynamical freezing effectively arrests the growth of chaos at special values of $A/\omega$.
Resource Availability
Lead Contact
Further information and requests should be directed to and will be fulfilled by Debanjan Chowdhury (debanjanchowdhury@cornell.edu)
Materials availability
This study did not generate new materials.
Data and code availability
  • All the theoretical data generated in this study has been deposited at Zenodo and is publicly available as of the date of publication.
  • The custom computer codes used to generate the results reported in this paper are available from the lead contact upon request.
Acknowledgments
@@
We thank S.D. Börner, A. Das, V. Fatemi, B. Kobrin, Y. Lensky, R. Moessner, E. Rosenberg, P. Roushan, S. Roy, V. Smelyanskiy, C. Tahan, G. Vidal and B. Ware for a number of useful discussions leading up to this work. RM thanks M. Bukov and S. Bandyopadhyay for many useful suggestions regarding QuSpin PYTHON package 77, 78, and thanks S.D. Börner for providing the Julia code dynamical.jl 68 for solving the semiclassical equations for the transmon arrays. RM is supported by a Fulbright-Nehru Grant No. 2877/FNDR/2023-2024 sponsored by the Bureau of Educational and Cultural Affairs of the United States Department of State. HG is supported by a Wilkins postdoctoral fellowship at Cornell University. KL acknowledges that this material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE – 2139899. DC is supported in part by a New Frontier Grant awarded by the College of Arts and Sciences at Cornell University and by a Sloan research fellowship from the Alfred P. Sloan foundation.
Author Contributions
This research was conceived and supervised by DC. RM performed all of the exact diagonalization computations. HG performed the Floquet-Magnus computations. KL performed the semiclassical computations. All the authors analyzed the results and wrote the paper.
Declaration of Interests
The authors declare no competing interests.
MAIN FIGURE TITLES AND LEGENDS
Figure 1. Schematic representation of a transmon array
The energy levels (red lines) for non-linear transmons with an unbounded spectrum coupled via capacitive interactions. The external Floquet drive, $f(t)$, couples capacitively to each transmon. $V$ is the capacitive coupling between the transmons.
Figure 2. Emergent conservation and entanglement entropy for the square-wave drive
In (a), (b), and (c) we show the stroboscopic time evolution of normalized $\widehat{N}(t)$ for different driving strengths for the square drive in Eq. (3a). The one-dimensional system size is $L=8$, with an on-site Hilbert space dimension $\mathcal{D}=3$. The drive frequency $\omega/(2\pi)=9.32$ GHz, and other parameters are as introduced in the main text. The initial state corresponds to the: (a) ground state, or thermal pure state with (b) $\beta^{-1}=0.79E_{J}$, (c) $\beta^{-1}=1.19E_{J}$, of $H(0)$. At freezing, there is a very slow decay of the normalized $\langle\widehat{N}(t)\rangle$ over time (inset). In (d), (e), and (f) we show the stroboscopic evolution of half-chain entanglement entropy at (red) and away from (blue) freezing for a square-wave drive for the same initial thermal states and same parameters as above. In the inset, we plot the relative change in entropy. At the freezing point, the entanglement growth is sufficiently suppressed, whereas, away from freezing, it increases rapidly before eventually saturating.
Figure 3. Renormalization of freezing points
(a) Time-averaged ${\cal{F}}$ (averaged over 100 Floquet cycles) as a function of $A/\omega$ for the cosine drive. We choose |$\ket{\psi(0)}$ from the eigenspectrum of $H(0)$. Primary freezing points, indicated by black arrows, shift from the zeroth-order prediction of freezing point in Eq. (40b) (dashed lines). The shift is due to the higher order corrections in the Magnus expansion, discussed in Sec. LABEL:results:FM. (b) The expectation value of the operator $\hat{N}$ is calculated with respect to the specific eigenstates of $H(0)$ and plotted against the corresponding eigen-energy. The states chosen to compute $\langle\widehat{N}(t)\rangle$ are indicated with the colored arrows and labeled $a,~{}b,~{}c$, respectively.
Figure 4. Emergent conservation and entanglement entropy for the cosine drive
In (a), (b), and (c) we show the stroboscopic time evolution of normalized $\widehat{N}(t)$ for the cosine drive given by Eq.(3b). The one-dimensional system size is $L=8$, with an on-site Hilbert space dimension $\mathcal{D}=3$. The drive frequency $\omega/(2\pi)=9.32$ GHz, and other parameters are as introduced in the main text. The initial state are chosen from the eigenstates of $H(0)$. The renormalized freezing points are obtained from Fig. LABEL:fig:cosonefidelity. At the renormalized freezing point, $\widehat{N}(t)$ becomes approximately conserved. Interestingly, in (b), the system rapidly loses its memory at $A/\omega$ = 2.4, the freezing point predicted by the zeroth-order Magnus expansion, that suggest significant contributions from the higher order corrections. In (d), (e), and (f) we show the stroboscopic evolution of half-chain entanglement entropy at (red), and away from (blue) freezing for a cosine wave drive for the same parameters and initial states, |$\ket{\psi(0)}$, as above. Even at the renormalized freezing point, we observe a significant growth in EE compared to the square wave drive, which can be attributed to higher-order corrections, even in the first-order Magnus expansion (see Sec. LABEL:results:FM).
Figure 5. Phenomenology of semi-classical freezing
The maximum Lyapunov exponent (MLE), $\lambda_{L}$, as a function of $A/\omega$ for a fixed drive frequency ( $\omega/(2\pi)=15$ GHz) with the square-wave drive, for a length $8$ one-dimensional chain initialized in the $1010$ state. The strong suppression of $\lambda_{L}$ coincides with the location of freezing points in Eq. 40a. Dashed horizontal line shows $\lambda_{L}$ for the static hamiltonian. (b) Correspondence between quantum and classical dynamics as a function of the ratio of drive frequency to transmon frequency (Defined in Eq. (4); see also Ref. 46 for a mapping between the transmon and Bose-Hubbard hamiltonian in the context of defining these frequencies) at freezing with $A/\omega=2$ kept fixed. The quantum expectation $\langle\widehat{N}(t)\rangle$ is computed by time-averaging over 100 Floquet cycles in the ground state of $H(0)$ which is plotted as the red curve. The blue curve is the corresponding maximum Lypunov exponent. For the quantum and semi-classical systems there exist similar thresholds for $\omega$ below which freezing no longer occurs. For both plots $L=8$, $E_{c}=330$ MHz, $E_{J}=12.58$ GHz , and $V=50$ MHz.
Figure 6. Phase space behavior of semi-classical freezing
Phase-space trajectories for two transmons (site $2$ and $5$, respectively) for a one-dimensional array with $L=8$ transmons, for initial state $1010$ and a square-wave drive with $\omega/(2\pi)=15$ GHz over $250$ cycles. The ratios of drive amplitude to frequency are (a) $A/\omega=7.3$, and (b) $A/\omega=8$ respectively. Note that away from freezing, the values of $n$ vary ergodically, while at freezing it remains tightly constrained around its initial value of 00.
Figure 7. Semi-classical freezing for the cosine drive
The maximum Lyapunov exponent as a function of $A/\omega$ for the cosine drive. The system with $L=8$ transmons is initialized in an $1010$ state with a drive frequency $\omega/(2\pi)=15$ GHz. The freezing points are indicated by arrows. The horizontal dashed line indicates $\lambda_{L}$ for the static Hamiltonian. The vertical dotted lines indicate the freezing points associated with the zeroth-order term in Eq. 40b.
Figure 8. Semi-classical freezing for IBM’s lattice geometries
(a) Heavy hexagon geometries used in IBM’s quantum processors. The light and dark blue qubits arranged in a checkerboard pattern are initialized as $1$ and 00, respectively. The colored regions show the geometries of IBM’s Falcon (green, $26$ qubits), Hummingbird (pink, $65$ qubits), Eagle (teal, $127$ qubits), and Osprey (orange, $433$ qubits) processors. (b) MLE as a function of $A/\omega$ for the different geometries with up to $433$ semi-classical qubits. All systems are driven at a fixed frequency ( $\omega/(2\pi)=15$ GHz) using the square-wave drive.
Figure 9. Dependency of semi-classical freezing on disorder
The maximum Lyapunov exponent as a function of disorder strength at the first and second renormalized freezing point, compared to the undriven case. Each curve is obtained for a one-dimensional array of length $8$ initialized in the $0101\ldots$ state and driven with a $\omega/(2\pi)=15$ GHz square-wave drive. Disorder is implemented by choosing $E_{J_{i}}\sim\mathcal{N}(12.58,\sigma^{2})$ GHz (normal distribution) independently for each site in the chain. Results are averaged over $1000$ disorder realizations. Typical disorder in IBM devices is on the order of $\sigma\approx 0.7$ GHz 62.
Figure 10. Dependency of freezing strength on drive frequency
(a) The time-averaged observable, ${\cal{F}}$ (Eq. (6)), as a function of $\widetilde{\omega}=\omega/\nu$ for $\nu$ given in Eq. (4). We choose |$\ket{\psi(0)}$ as the ground state of $H(0)$ and fix $A/\omega=2.4048$ at the first freezing point within the zeroth-order Magnus expansion. The scaling is predominantly proportional to $1/\omega$ (Eq. (14)). (b) ${\cal{F}}$ at the first renormalized freezing value of $A/\omega$ (taking into account the zeroth and first-order corrections). The other parameters are the same as in (a).
Supplementary information for “Arresting Quantum Chaos Dynamically in Transmon Arrays”
Appendix A Freezing for Large Hilbert-Space Dimension
Here we present additional results for larger local Hilbert space dimensions, in order to fully capture the relevant freezing phenomenology. To account for the chaotic effects of the driven transmon, we must truncate the local Hilbert space such that all of the energy levels extending to the top of the cosine potential are included. The approximate eigen-energy of the single transmon is given by
$\epsilon_{n}=\sqrt{8E_{J}E_{c}}(n+\frac{1}{2})-\frac{E_{c}}{12}(6n^{2}+6n+3).$ (S.1)
For our choice of parameters, $E_{J}=12.58$ GHz and $E_{c}=330$ MHz, the number of bound states corresponding to the maximum number $n_{\text{max}}$ such that the $\epsilon_{n_{\text{max}}}$ remains below the barrier height, is approximately given by $E_{J}$. The number of bound state in our case is thus $N_{\text{bound}}\approx 6$. Results for local Hilbert space dimension ${\cal{H}}_{D}=8$ for a one-dimensional chain of $4$ transmons are shown in Fig. LABEL:fig:firstorderscaling. To correctly approximate the cosine potential, we expand the potential to $8^{\text{th}}$ order in $\phi$ as
$\cos{(\hat{\phi})}=I-\frac{\hat{\phi}^{2}}{2!}+\frac{\hat{\phi}^{4}}{4!}-\frac{\hat{\phi}^{6}}{6!}+\frac{\hat{\phi}^{8}}{8!}.$ (S.2)
We find that freezing still occurs near the $A/\omega=2$ zeroth-order freezing point.
In Fig.LABEL:fig:HilbertspaceSystemsize, we systematically study the freezing phenomenology as a function of both system size and onsite Hilbert space dimension for the square wave drive. We find that the freezing point is robust. In (a), we observe that the expectation value of $N(t)/N(0)$ converges as a function of system size when the onsite Hilbert space dimension is kept fixed. On the other hand, when the Hilbert space dimension is increased while keeping the system size fixed (as shown in Fig. LABEL:fig:HilbertspaceSystemsize (b)), the freezing point is slightly shifted away from the predictions of the leading-order Magnus expansion.
Away from the freezing point, there exist variations of $N(t)/N(0)$ concerning the local Hilbert space dimension $\mathcal{D}$. This phenomenon is anticipated since, away from the freezing point, the system rapidly thermalizes and enters a chaotic regime, as evidenced by the entanglement entropy depicted in Fig. LABEL:fig:ndtstrobo(d),(e),(f) in the main text. The saturation value of the entanglement entropy is expected to sensitively depend on $\mathcal{D}$, and similar occurrences are expected in $N(t)/N(0)$. However, at the freezing point, due to the emergent conservation and integrability, the state remains localized in the Fock basis, and is therefore independent of $\mathcal{D}$.
Appendix B Analytical details of the Floquet-Magnus Expansion
@@
Here we provide additional details regarding the Floquet-Magnus expansion.
B.1 General Expression for the Floquet Hamiltonian
@@
$U_{I}(t)={\mathcal{T}}\exp\left(-i\int_{0}^{t}{\rm d}t^{\prime}\mathcal{H}_{\text{mov}}(t^{\prime})\right)\,,$ (S.3)
@@
$i\partial_{t}U_{I}(t)=\mathcal{H}_{\text{mov}}(t)U_{I}(t)\,.$ (S.4)
@@
$\begin{split}\partial_{t}e^{\Omega(t)}&=\int_{0}^{1}{\rm d}\lambda e^{\lambda\Omega(t)}\partial_{t}\Omega(t)e^{(1-\lambda)\Omega(t)}\\ &=h({\rm ad}_{\Omega(t)})\partial_{t}\Omega e^{\Omega(t)}\,,\end{split}$ (S.5)
where $h(x)=(e^{x}-1)/x$ and ${\rm ad}_{X}Y=[X,Y]$. Combining this with Eq. (S.4), we obtain,
$\partial_{t}\Omega(t)=-ih\left({\rm ad}_{\Omega(t)}\right)\mathcal{H}_{\text{mov}}(t)\,.$ (S.6)
@@
$\mathcal{H}_{\text{eff}}^{0}$ $=\frac{1}{T}\int_{0}^{T}{\rm d}t\mathcal{H}_{\text{mov}}(t)\,,$ (S.7)
$\mathcal{H}_{\text{eff}}^{1}$ $=\frac{1}{2iT}\int_{0}^{T}{\rm d}t_{1}\int_{0}^{t_{1}}{\rm d}t_{2}\left[\mathcal{H}_{\text{mov}}(t_{1}),\mathcal{H}_{\text{mov}}(t_{2})\right]\,,$ (S.8)
$\mathcal{H}_{\text{eff}}^{2}$ $=\frac{-1}{6T}\int_{0}^{T}{\rm d}t_{1}\int_{0}^{t_{1}}{\rm d}t_{2}\int_{0}^{t_{2}}{\rm d}t_{3}$ (S.9)
@@
B.2 Magnus Expansion of the Transmon Hamiltonian
In this section we explicitly evaluate Eqs. (S.7)-(S.9) for the transmon Hamiltonian. The co-moving Hamiltonian is given by
$\mathcal{H}_{\text{mov}}(t)=\mathcal{H}^{0}_{\widehat{n}}-E_{J}\sum_{k}\cos({\widehat{\varphi}}_{k}-\theta_{k}(t))\,,$ (S.10)